OpenStudy (anonymous):

How do i work this out i am lost on how to do it? If an object moves in uniform circular motion in a circle of radius R = 1.0 meter, and the object takes 4.0 seconds to complete ten revolutions, calculate the centripetal acceleration.

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Peroidic time T = 0.4 s ω = 2π/T ω² = 4π²/0.4² = 246.7 (rad/s)² centripetal accel. = Rω² = 1.0*246.7 .. .. ►246.7 m/s²

OpenStudy (yanasidlinskiy):

A different way... 10 revolutions / 4 secs = 2.5 rps 2.5 x 2pi = 15.708 rad/s omega x radius = tangential velocity m/s 15.708 x 1 = 15.708 m/s centripetal acceleration = v^2/r 15.708^2/1 = 246.74 m/s^2

OpenStudy (anonymous):

Okay thank you!

OpenStudy (yanasidlinskiy):

Your Welcome :D